OpenStudy (anonymous):

6-2(square root)3n = 0

OpenStudy (lgbasallote):

\[\large 6 - 2\sqrt{3n} = 0?\]

OpenStudy (anonymous):

yaa

OpenStudy (lgbasallote):

okay...step 1: add \(2\sqrt{3n}\) to both sides to make it positive

OpenStudy (anonymous):

step two?

OpenStudy (lgbasallote):

divide both sides by 2

OpenStudy (lgbasallote):

so it's just \(\sqrt{3n}\)on one side

OpenStudy (anonymous):

then its \[3 = \sqrt{3n}\] right?

OpenStudy (lgbasallote):

yep! square both sides to get n out of the root

OpenStudy (anonymous):

show meeee

OpenStudy (lgbasallote):

\[(3)^2 = (\sqrt{3n})^2\] the square root will cancel \[9 = 3n\] got it?

OpenStudy (anonymous):

yeah ! Thank youuu

OpenStudy (lgbasallote):

last step is to divide by 3 :D

OpenStudy (anonymous):

i know that :) lol

OpenStudy (lgbasallote):

good ^_^